Two cylindrical drums have their volumes in ratio 8:27


1
Expert's answer
2022-01-26T16:01:49-0500
"V=\pi r^2 h"

Suppose that these two cylinders are similar


"r_1\/r_2=h_1\/h_2"

Then


"\dfrac{V_1}{V_2}=\dfrac{\pi r_1^2h_1 }{\pi r_2^2h_2}=(\dfrac{r_1}{r_2})^3=\dfrac{8}{27}"

"\dfrac{r_1}{r_2}=\dfrac{2}{3}=\dfrac{h_1}{h_2}"